# Category $\mathcal{O}$ for Lie superalgebras

### v1

@@ 3.1. The Definition
Let $C_{(\mathfrak{g},\mathfrak{a})}$ be the full subcategory of $\mathfrak{g}$-Mod consisting of finitely $\mathfrak{a}$-semisimple $\mathfrak{g}$-modules. That is, each module $M\in C_{(\mathfrak{g},\mathfrak{a})}$ satisfies the following conditions:
@@
    $M$ is a direct sum of finite-dimensional $\mathfrak{a}$-modules.
@@
Assume that $\mathfrak{g}$ admits a decomposition $\mathfrak{g}=\mathfrak{n}^{+}\oplus\mathfrak{a}\oplus\mathfrak{n}^{-}$ of Lie sub-superalgebras. We define a new Category $\mathcal{O}$ as follows. Let $\mathcal{O}^{\mathfrak{p}}$ be the full subcategory of $U(\mathfrak{g})$-Mod such that each module $M\in\mathcal{O}^{\mathfrak{p}}$ satisfies the following conditions:
@@ 3.3. Important facts about ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
    The composition multiplicities $[Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ of simple modules in Verma modules are universally bounded in ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$.
@@
Part (a) and (b) are standard facts. For part (c), let $\gamma\in\Lambda_{{\bar{0}}}$. Then we have $[Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]\neq 0$ for only finitely many $\tau\in\Lambda_{{\bar{0}}}$ because every $\tau\in\Lambda_{{\bar{0}}}$ for which $L_{{\bar{0}}}(\tau)$ appears as a composition factor of $Z_{{\bar{0}}}(\gamma)$ must be in the finite $W_{{\mathfrak{p}}_{{\bar{0}}}}$-orbit of $\gamma$ under the dot-action. Since ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$ has finite length, we also have $[Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]<\infty$. Since there are finitely many blocks (up to Morita equivalence), we can take the maximum of these non-zero numbers to get a bound. ∎
@@ 4.1. Category $\mathcal{O}$ for Semisimple Lie Algebras
In this context, Theorem A translates into the isomorphisms $\mathrm{Ext}_{\mathcal{O}^{\mathfrak{p}}}^{i}(M,N)\cong\mathrm{Ext}^{i}_{(\mathfrak{g},\mathfrak{a})}(M,N)$ for all $i\geq 0$, and $M,N\in\mathcal{O}^{\mathfrak{p}}$. The isomorphisms for the special case where $\mathfrak{a}=\mathfrak{h}$ are well-known. It was stated in [De80] (without proof), and has appeared an an exercise in Kumar’s book [Ku02, Exercise 3.3.E(1)]. A proof for this special case can be found in [CM15, Theorem 9], in which the statement is equivalent to that the inclusion functor $\mathcal{O}\hookrightarrow\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ is extension full.
@@ 4.3. Quasi-reductive Lie Superalgebras
Suppose that $\mathfrak{g}$ is a quasi-reductive (not necessarily simple) Lie superalgebra. Mazorchuk’s parabolic Category $\mathcal{O}$ [Maz14] is an example of our Category $\mathcal{O}^{\mathfrak{p}}$, where $\mathfrak{p}$ is a principal parabolic subalgebra in the sense of Grantcharov and Yakimov [GY13]. In particular, such a Category $\mathcal{O}^{\mathfrak{p}}$ recovers Brundan’s Category $\mathcal{O}_{m|n}$ [Br03] for $\mathfrak{gl}(m|n)$ and Category $\mathcal{O}_{n}$ [Br04] for $\mathfrak{q}(n)$, as well as the parabolic categories over $\mathfrak{gl}(m|n)$ and over $\mathfrak{osp}(m|n)$ which appeared in the super duality [CLW15] established by Cheng, Lam and Wang.
@@ 5.1. Relative cohomology for rings
    We say this sequence is $(R,S)$-exact if it is exact as a sequence of $R$-supermodules and if, when viewed as a sequence of $S$-supermodules, Ker $\text{Ker }f_{i}$ is a direct summand of $M_{i}$ for all $i.$
@@
An equivalent definition of an $R$-module, $M$, being $(R,S)$-projective is that $M$ is a direct summand of $R\otimes_{S}N$ for some $S$-module $N$. From this it is clear that any projective $R$-module is $(R,S)$-projective. Furthermore, any $R$-supermodule $M$ admits an $(R,S)$-projective resolution. We can define cohomology as follows: let $P_{\bullet}\to M$ be a $(R,S)$-projective resolution and let
@@
For shorthand notation, $({\mathfrak{g}},{\mathfrak{h}})$-injective will mean $(U({\mathfrak{g}}),U({\mathfrak{h}}))$-injective. Note that if $N$ is $({\mathfrak{g}},{\mathfrak{h}})$-injective then $\text{Hom}_{\mathbb{C}}(M,N)$ is $({\mathfrak{g}},{\mathfrak{h}})$-injective. Also, $F$ is $({\mathfrak{g}},{\mathfrak{h}})$-split. These results use the fact that $F(-)$ is exact because one is taking homomorphisms over the field ${\mathbb{C}}$.
@@ 5.3. Explicit complex for relative Lie superalgebra cohomology
In this section we define (relative) Lie superalgebra cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ where $\mathfrak{g}$ is a Lie superalgebra and $\mathfrak{h}\subseteq\mathfrak{g}$ is a Lie subsuperalgebra. This will be defined via an explicit complex which enables us to make explicit computations. Later this will be related to the relative cohomology for a pair of superalgebras.
@@
(a) First note that the second sum of (5.3.3) are all zero when $M\cong{\mathbb{C}}$. Next observe that $[\mathfrak{g}_{{\bar{1}}},\mathfrak{g}_{{\bar{1}}}]\subseteq\mathfrak{g}_{{\bar{0}}}$, thus each $[x_{i},x_{j}]$ is always zero $\mathfrak{g}/\mathfrak{g}_{{\bar{0}}}$, and the first sum of (5.3.3) is zero. Hence, differentials $d^{p}$ are all zero, and thus the cohomology ring identifies with the cochains.
@@ 5.4. Relative Lie Superalgebra Cohomology: Spectral Sequences
$E_{1}^{p,q}={\operatorname{H}}^{q}(\mathfrak{h},{\mathfrak{q}},\mathrm{Hom}_{\mathbb{C}}(\Lambda^{p}(\mathfrak{g}/\mathfrak{h}),M))\Rightarrow{\operatorname{H}}^{q}(\mathfrak{g},\mathfrak{h},M).$
@@
Let ${\mathfrak{q}}\leq\mathfrak{g}_{{\bar{0}}}\leq\mathfrak{g}$ be inclusions of finite-dimensional Lie superalgebras, and let $M$ be a $\mathfrak{g}$-module. Then there exists a first quadrant spectral sequence
@@ 5.5. Cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ and finite generation
We will always assume $\mathfrak{g}$ is a classical Lie superalgebra. A $\mathfrak{g}$-supermodule will always be assumed to be an object in the category $\mathcal{C}=\mathcal{C}_{(\mathfrak{g},\mathfrak{g}_{{\bar{0}}})}$ and a finite dimensional $\mathfrak{g}$-supermodule will always mean an object in the category $\mathcal{F}=\mathcal{F}_{(\mathfrak{g},\mathfrak{g}_{{\bar{0}}})}.$
@@
    There is a spectral sequence $\{E_{r}^{p,q}\}$ which computes cohomology and satisfies
@@ 6.3. A Spectral Sequence
Assume that $\operatorname{Ext}^{1}_{U({\mathfrak{a}}_{{\bar{0}}})}(M,N)=0$ for all finite-dimensional $U({\mathfrak{a}}_{{\bar{0}}})$-modules. Then the functor ${\mathcal{F}}$ takes relative injective objects in $\mathcal{C}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}$ to injective objects in $\mathcal{O}^{\mathfrak{p}}$.
@@
where $M\in\mathcal{C}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}$ and $N\in\mathcal{O}^{\mathfrak{p}}$.
@@
Let $\Lambda$ (resp. $\Lambda_{{\bar{0}}}$) with the indexing of the corresponding simple and projective modules be as in Section 6.2. For $\lambda\in\Lambda$, $j\geq 0$ and $N\in\mathcal{O}^{\mathfrak{p}}$,
$[R^{j}\mathcal{F}(N):L(\lambda)]=\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}(P(\lambda),N).$ (6.3.5)
@@ 6.5. Proof of Theorem B
with $M^{*}\otimes N$ in $\mathcal{O}^{\mathfrak{p}}$. According to Theorem 5.5.1(c), it suffices to prove that $C^{\bullet}({\mathfrak{g}},{\mathfrak{g}}_{{\bar{0}}},\widehat{N})$ is a finitely generated module over $R=S^{\bullet}({\mathfrak{g}}_{{\bar{1}}}^{*})^{G_{{\bar{0}}}}$ where $\widehat{N}$ is in $\mathcal{O}^{\mathfrak{p}}$. For ${\mathfrak{g}}$ classical simple, it was shown in [BKN10a, Theorem 3.4] that as an $R$-module one has a harmonic decomposition
@@
$C^{\bullet}({\mathfrak{g}},{\mathfrak{g}}_{{\bar{0}}},M)\cong\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}(S^{\bullet}({\mathfrak{g}}_{{\bar{1}}}),\widehat{N})\cong S^{\bullet}({\mathfrak{g}}_{{\bar{1}}}^{*})^{G_{{\bar{0}}}}\otimes\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}({\mathbb{C}},[\text{ind}_{H}^{G_{{\bar{0}}}}{\mathbb{C}}]_{\bullet}\otimes\widehat{N}).$ (6.5.4)
It suffices to show that $\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}({\mathbb{C}},[\text{ind}_{H}^{G_{{\bar{0}}}}{\mathbb{C}}]\otimes\widehat{N})$ is finite-dimensional.
@@
$\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}({\mathbb{C}},[\text{ind}_{H}^{G_{{\bar{0}}}}{\mathbb{C}}]\otimes\widehat{N})\hookrightarrow\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}({\mathbb{C}},[\text{Hom}_{U({\mathfrak{h}})}(U({\mathfrak{g}}),{\mathbb{C}})]\otimes\widehat{N}).$ (6.5.5)
@@
$\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}({\mathbb{C}},[\text{Hom}_{U({\mathfrak{h}})}(U({\mathfrak{g}}),{\mathbb{C}})]\otimes\widehat{N})\cong\text{Hom}_{U({\mathfrak{h}})}({\mathbb{C}},\widehat{N}).$ (6.5.6)
@@
$\text{Hom}_{U({\mathfrak{h}})}({\mathbb{C}},\widehat{N})\hookrightarrow\text{Hom}_{U(\hat{\mathfrak{t}})}({\mathbb{C}},\widehat{N})\cong\text{Hom}_{\widehat{T}}({\mathbb{C}},\widehat{N}).$ (6.5.7)
With this series of injective maps, one is reduced to showing that $\text{Hom}_{\widehat{T}}({\mathbb{C}},\widehat{N})$ is finite-dimensional.
@@ 7.1. Ext Orthogonality
Let $\Lambda_{\mathfrak{a}}$ index the non-isomorphic simple $\mathfrak{a}$-modules in $\mathcal{C}_{(\mathfrak{a},\mathfrak{a}_{\bar{0}})}$. For $\lambda_{0}\in\Lambda_{\mathfrak{a}}$, let $L(\lambda_{0})$ be the corresponding simple module and $I_{\mathfrak{a}}(\lambda_{0})$ be its injective hull. Any simple module in $\mathcal{O}^{\mathfrak{p}}$ can be realized as a quotient of $U({\mathfrak{g}})\otimes_{U({\mathfrak{p}})}L_{\mathfrak{a}}(\lambda_{0})$ for some $\lambda_{0}\in\Lambda_{\mathfrak{a}}$. One can now use the argument in [HN91, Section 3] to show that the simples in $\mathcal{O}^{\mathfrak{p}}$ are also indexed by $\Lambda_{{\mathfrak{a}}}$ (before considering parity). Set $\Lambda:=\Lambda_{\mathfrak{a}}\times\{0,1\}$. For any $\lambda=(\lambda_{0},i)\in\Lambda$, let $L(\lambda)$ be the corresponding simple module in $\mathcal{O}^{\mathfrak{p}}$. Here it is important to remember that $\lambda_{0}$ and $\lambda$ should not be regarded as weights.
@@
$E_{2}^{i,j}=\mathrm{Ext}^{i}_{(\mathfrak{a},\mathfrak{a}_{\bar{0}})}(\mathbb{C},\mathrm{Ext}^{j}_{U(\mathfrak{n}^{-})}(N,\mathbb{C})\otimes I_{\mathfrak{a}}(\mu_{0}))\Rightarrow\mathrm{Ext}^{i+j}_{(\mathfrak{p}^{-},\mathfrak{a}_{\bar{0}})}(N,I_{\mathfrak{a}}(\mu_{0}).$ (7.1.9)
@@
For any $\nu\leq\mu$, by (c) we have $\mathrm{Ext}^{1}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V)=0$, and thus $\nabla_{\mathrm{inj}}(\mu)\hookrightarrow V$ by Lemma 7.2.1, i.e., $V$ has a submodule $V^{\prime}$ that is isomorphic to $\nabla_{\mathrm{inj}}(\mu)$. By Proposition 7.1.1, $V^{\prime}$ satisfies (b). Also, $V^{\prime\prime}:=V/V^{\prime}$ satisfies (c). Note that $\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V^{\prime\prime})\cong\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V)$ for all $\nu\neq\mu$ whereas $\dim\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\mu),V^{\prime\prime})=\dim\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\mu),V)-1$. We can then iterate this construction to obtain a $\nabla_{\mathrm{inj}}$-filtration for $V$. ∎

### v2

@@ 3.1. The Definition
Let $\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ be the full subcategory of $\mathfrak{g}$-Mod with the following properties.
@@
    Every module $M$ is a direct sum of finite-dimensional $\mathfrak{a}$-modules.
@@
Assume that $\mathfrak{g}$ admits a decomposition $\mathfrak{g}=\mathfrak{n}^{+}\oplus\mathfrak{a}\oplus\mathfrak{n}^{-}$ of Lie sub-superalgebras. We define a new Category $\mathcal{O}$ as follows. Let $\mathcal{O}^{\mathfrak{p}}$ be the full subcategory of $\mathfrak{g}$-Mod such that each module $M\in\mathcal{O}^{\mathfrak{p}}$ satisfies the following conditions:
@@ 3.3. Important facts about ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
    The composition multiplicities $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ of simple modules in Verma modules are universally bounded in ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$.
@@
Part (a) and (b) are standard facts. For part (c), let $\gamma\in\Lambda_{{\bar{0}}}$. Then we have $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]\neq 0$ for only finitely many $\tau\in\Lambda_{{\bar{0}}}$ because every $\tau\in\Lambda_{{\bar{0}}}$ for which $L_{{\bar{0}}}(\tau)$ appears as a composition factor of $Z_{{\bar{0}}}(\gamma)$ must be in the finite $W_{{\mathfrak{p}}_{{\bar{0}}}}$-orbit of $\gamma$ under the dot-action. Since ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$ has finite length, we also have $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]<\infty$. Since there are finitely many blocks (up to Morita equivalence), we can take the maximum of these non-zero numbers to get a bound. ∎
@@ 4.1. Category $\mathcal{O}$ for Semisimple Lie Algebras
In this context, Theorem A translates into the isomorphisms $\mathrm{Ext}_{\mathcal{O}^{\mathfrak{p}}}^{i}(M,N)\cong\mathrm{Ext}^{i}_{(\mathfrak{g},\mathfrak{a})}(M,N)$ for all $i\geq 0$, and $M,N\in\mathcal{O}^{\mathfrak{p}}$. The isomorphisms for the special case where $\mathfrak{a}=\mathfrak{h}$ are well-known. It was stated in [De80] (without proof), and has appeared as an exercise in Kumar’s book [Ku02, Exercise 3.3.E(1)]. A proof for this special case can be found in [CM15, Theorem 9], in which the statement is equivalent to that the inclusion functor $\mathcal{O}\hookrightarrow\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ is extension full.
@@ 4.3. Quasi-reductive Lie Superalgebras
Suppose that $\mathfrak{g}$ is a quasi-reductive (not necessarily simple) Lie superalgebra. Mazorchuk’s parabolic Category $\mathcal{O}$ [Maz14] is an example of our Category $\mathcal{O}^{\mathfrak{p}}$, where $\mathfrak{p}$ is a principal parabolic subalgebra as defined by Grantcharov and Yakimov [GY13]. In particular, such a Category $\mathcal{O}^{\mathfrak{p}}$ recovers Brundan’s Category $\mathcal{O}_{m|n}$ [Br03] for $\mathfrak{gl}(m|n)$ and Category $\mathcal{O}_{n}$ [Br04] for $\mathfrak{q}(n)$, as well as the parabolic categories over $\mathfrak{gl}(m|n)$ and over $\mathfrak{osp}(m|n)$ which appeared in the super duality [CLW15] established by Cheng, Lam and Wang.
@@ 5.1. Relative cohomology for rings
    We say this sequence is $(R,S)$-exact if it is exact as a sequence of $R$-supermodules and if, when viewed as a sequence of $S$-supermodules, $\text{Ker }f_{i}$ is a direct summand of $M_{i}$ for all $i.$
@@
An equivalent definition of an $R$-module, $M$, being $(R,S)$-projective is that $M$ is a direct summand of $R\otimes_{S}N$ for some $S$-module NN. From this it is clear that any projective $R$-module is $(R,S)$-projective. Furthermore, any $R$-supermodule $M$ admits an $(R,S)$-projective resolution. We can define cohomology as follows: let $P_{\bullet}\to M$ be a $(R,S)$-projective resolution and let
@@
From this point on, we write $\mathrm{Ext}^{\bullet}_{(\mathfrak{g},\mathfrak{h})}(M,N)$ to denote $\mathrm{Ext}^{\bullet}_{({U}(\mathfrak{g}),{U}(\mathfrak{h}))}(M,N)$, under the conditions of the theorem.
@@
For shorthand notation, $({\mathfrak{g}},{\mathfrak{h}})$-injective will mean $(U({\mathfrak{g}}),U({\mathfrak{h}}))$-injective. Note that if $N$ is $({\mathfrak{g}},{\mathfrak{h}})$-injective then $\text{Hom}_{\mathbb{C}}(M,N)$ is $({\mathfrak{g}},{\mathfrak{h}})$-injective. Also, $F$ is $({\mathfrak{g}},{\mathfrak{h}})$-split. These results use the fact that F()F(-) is exact because one is taking homomorphisms over the field $\mathbb{C}$}.
@@ 5.3. Explicit complex for relative Lie superalgebra cohomology
In this section we define (relative) Lie superalgebra cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ where $\mathfrak{g}$ is a Lie superalgebra and $\mathfrak{h}\subseteq\mathfrak{g}$ is a Lie subsuperalgebra. This will be defined via a complex which enables us to make explicit computations. Later this will be related to the relative cohomology for a pair of superalgebras.
@@
(a) First note that the second sum of (5.3.3) are all zero when $M\cong{\mathbb{C}}$. Next observe that $\mathfrak{g}_{{\bar{1}}},\mathfrak{g}_{{\bar{1}}}]\subseteq\mathfrak{g}_{{\bar{0}}}$, thus each $x_{i},x_{j}]$ is always zero $\mathfrak{g}/\mathfrak{g}_{{\bar{0}}}$, and the first sum of (5.3.3) is zero. Hence, differentials $d^{p}$ are all zero, and thus the cohomology ring identifies with the cochains.
@@ 5.4. Relative Lie Superalgebra Cohomology: Spectral Sequences
$E_{1}^{p,q}={\operatorname{H}}^{q}(\mathfrak{h},{\mathfrak{q}},\mathrm{Hom}_{\mathbb{C}}(\Lambda^{p}(\mathfrak{g}/\mathfrak{h}),M))\Rightarrow{\operatorname{H}}^{q}(\mathfrak{g},\mathfrak{q},M).$
@@
Let ${\mathfrak{q}}\leq\mathfrak{g}_{{\bar{0}}}\leq\mathfrak{g}$ be inclusions of finite-dimensional Lie superalgebras, and let $M$ be a finite-dimensional $\mathfrak{g}$-module. Then there exists a first quadrant spectral sequence
@@ 5.5. Cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ and finite generation
We will always assume $\mathfrak{g}$ is a classical (quasi-reductive) Lie superalgebra. A $\mathfrak{g}$-supermodule will always be assumed to be an object in the category $\mathcal{C}=\mathcal{C}_{(\mathfrak{g},\mathfrak{g}_{{\bar{0}}})}$ and a finite dimensional $\mathfrak{g}$-supermodule will always mean an object in the category $\mathcal{F}=\mathcal{F}_{(\mathfrak{g},\mathfrak{g}_{{\bar{0}}})}.$
@@
    There is a spectral sequence $\{E_{r}^{p,q}\}$ when $M$ is finite-dimensional which computes cohomology and satisfies
@@ 6.3. A Spectral Sequence
Assume that for all finite-dimensional $U({\mathfrak{a}}_{{\bar{0}}})$-modules $M$ and $N$, $\operatorname{Ext}^{1}_{U({\mathfrak{a}}_{{\bar{0}}})}(M,N)=0$. Then the functor $\mathcal{F}$} takes relative injective objects in $\mathcal{C}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}$ to injective objects in $\mathcal{O}^{\mathfrak{p}}$.
@@
where $M,N\in\mathcal{C}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}$.
@@
Let $\Lambda$ (resp. $\Lambda_{{\bar{0}}}$) with the indexing of the corresponding simple and projective modules be as in Section 6.2. For $\lambda\in\Lambda$, $j\geq 0$ and $N\in\mathcal{C}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}$,
$R^{j}\mathcal{F}(N):L(\lambda)]=\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{a}_{\bar{0}})}(P(\lambda),N).$ (6.3.5)
@@ 6.5. Proof of Theorem B
[…]
@@
[…]
Set ${\mathfrak{h}}=\text{Lie }H$. Observe that $\text{ind}_{H}^{G_{{\bar{0}}}}{\mathbb{C}}$ is locally finite induction, so there exists an injective homomorphism $\text{ind}_{H}^{G_{{\bar{0}}}}{\mathbb{C}}\hookrightarrow\text{Hom}_{U({\mathfrak{h}})}(U({\mathfrak{g}}),{\mathbb{C}})$. This induces an injective map
[…]
@@
[…]
@@
[…]
@@ 7.1. Ext Orthogonality
Let $\Lambda_{\mathfrak{a}}$ index the non-isomorphic simple $\mathfrak{a}$-modules in $\mathcal{C}_{(\mathfrak{a},\mathfrak{a}_{\bar{0}})}$. For $\lambda_{0}\in\Lambda_{\mathfrak{a}}$, let $L_{\mathfrak{a}}(\lambda_{0})$ be the corresponding simple module and $I_{\mathfrak{a}}(\lambda_{0})$ be its injective hull. Any simple module in $\mathcal{O}^{\mathfrak{p}}$ can be realized as a quotient of $U({\mathfrak{g}})\otimes_{U({\mathfrak{p}})}L_{\mathfrak{a}}(\lambda_{0})$ for some $\lambda_{0}\in\Lambda_{\mathfrak{a}}$. One can now use the argument in [HN91, Section 3] to show that the simples in $\mathcal{O}^{\mathfrak{p}}$ are also indexed by $\Lambda_{{\mathfrak{a}}}$ (before considering parity). Set $\Lambda:=\Lambda_{\mathfrak{a}}\times\{0,1\}$. For any $\lambda=(\lambda_{0},i)\in\Lambda$, let $L(\lambda)$ be the corresponding simple module in $\mathcal{O}^{\mathfrak{p}}$. Here it is important to remember that $\lambda_{0}$ and $\lambda$ should not be regarded as weights.
@@
$E_{2}^{i,j}=\mathrm{Ext}^{i}_{(\mathfrak{a},\mathfrak{a}_{\bar{0}})}(\mathbb{C},\mathrm{Ext}^{j}_{U(\mathfrak{n}^{-})}(N,\mathbb{C})\otimes I_{\mathfrak{a}}(\mu_{0}))\Rightarrow\mathrm{Ext}^{i+j}_{(\mathfrak{p}^{-},\mathfrak{a}_{\bar{0}})}(N,I_{\mathfrak{a}}(\mu_{0})).$ (7.1.9)
@@
Furthermore, if $I(\lambda)$ is the injective hull of $L(\lambda)$ in ${\mathcal{O}}^{\mathfrak{p}}$, then one has the following reciprocity law:
$I(\lambda):\nabla_{\mathrm{inj}}(\gamma)]=[\Delta_{\mathrm{irr}}(\gamma):L(\lambda)]$
for $\lambda,\gamma\in\Lambda$.
@@
For any $\nu\leq\mu$, by (c) we have $\mathrm{Ext}^{1}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V)=0$, and thus $\nabla_{\mathrm{inj}}(\mu)\hookrightarrow V$ by Lemma 7.2.1, i.e., $V$ has a submodule $V^{\prime}$ that is isomorphic to $\nabla_{\mathrm{inj}}(\mu)$. By Proposition 7.1.1, $V^{\prime}$ satisfies (b). Also, $V^{\prime\prime}:=V/V^{\prime}$ satisfies (c). Note that $\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V^{\prime\prime})\cong\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\nu),V)$ for all $\nu\neq\mu$ whereas $\dim\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\mu),V^{\prime\prime})=\dim\mathrm{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\mu),V)-1$. We can then iterate this construction to obtain a $\nabla_{\mathrm{inj}}$-filtration for $V$.
Finally, observe that by (c), $I(\lambda)$ has a $\nabla_{\mathrm{inj}}$-filtration, and one can count the number of times such a factor occurs by apply $\text{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\gamma),-)$ to short exact sequences of modules with a $\nabla_{\mathrm{inj}}$-filtration. Therefore,
$I(\lambda):\nabla_{\mathrm{inj}}(\gamma)]=\dim\text{Hom}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\gamma),I(\lambda))=[\Delta_{\mathrm{irr}}(\gamma):L(\lambda)].$